The boiling point of an aqueous 1.83 m (NH4)2SO4 (molar mass = 132.15 g/mol) solution is 102.5°C. Determine the value of the van't Hoff factor for this solute if the Kb for water is 0.512°C/m.

Answer's 2.7 but I have no idea how to solve for this, can someone show me how to solve with LOTS of steps? Thanks.
Answer #1
Basically you're working with the equation ?T(boiling) = Kb * m * i and you want to solve for i.

The normal boiling point for water is 100 C so the the ?Tb = (102.5 - 100 = 2.5). From here you just have to plug everything in. The solution is 1.83 molal, and the Kb is 0.512. So i = ?T/(kb * m) = 2.5/(.512 * 1.83) = 2.67